OpenStudy (anonymous):

Seven subtracted from 3 times a number is equal to 9 more than 5 times the number. What is the number?

OpenStudy (midhun.madhu1987):

Let the number be x. Then, 3 times a number = 3x 7 subtracted from it = (3x - 7) 3x - 7 = 5x + 9 Solve for x.

OpenStudy (anonymous):

how

OpenStudy (midhun.madhu1987):

Move all terms containing x to the left side and the numbers to right side..

OpenStudy (anonymous):

so 3x + 5x = -7 + 9?

OpenStudy (anonymous):

8x = 2

OpenStudy (midhun.madhu1987):

+ becomes - when it moves across the equal sign

OpenStudy (midhun.madhu1987):

and - bcomes +

OpenStudy (anonymous):

3x - 5x = 7 + 9

OpenStudy (anonymous):

-2x = 16

OpenStudy (midhun.madhu1987):

yes... perfect.. so x = -8 :)

OpenStudy (anonymous):

thanks :)

OpenStudy (midhun.madhu1987):

yw :)